OpenStudy (bebe):

what is 3/4 subtracted by 1/4

OpenStudy (bebe):

how would i find the sum exactly

OpenStudy (akshay_budhkar):

you want the sum or you want to subtract?

OpenStudy (bebe):

the sum

OpenStudy (akshay_budhkar):

the sum is 3/4+1/4=4/4=1

OpenStudy (bebe):

thats the answer

OpenStudy (akshay_budhkar):

cool...... but you wrote in question you want to subtract

OpenStudy (bebe):

i said find the sum

OpenStudy (akshay_budhkar):

you get ur answer right? u get how to solve it right?

OpenStudy (anonymous):

1/4 - 3/4 cannot subtract. But 3/4 + 1/4 = 4/4 =1 whole